### v1

# A System Level Approach to LQR Control of the Diffusion Equation

Abstract
The continuous-time, infinite horizon LQR problem for the diffusion equation over the unit circle with fully distributed actuation is considered. It is well-known that the solution to this problem can be obtained from the solution to an operator-valued algebraic Riccati equation. Here, it is demonstrated that this solution can be equivalently obtained by solving an $H_{2}$ control problem through a closed-loop design procedure that is analogous to the “System Level Synthesis” methodology previously developed for systems over a discrete spatial domain and/or over a finite time horizon. The presented extension to the continuous spatial domain and continuous and infinite-horizon time setting admits analytical solutions that may complement computational approaches for discrete or finite-horizon settings. It is further illustrated that spatio-temporal constraints on the closed-loop responses can be incorporated into this new formulation in a convex manner.
I Introduction
As large-scale spatially-distributed systems with immense amounts of sensing and actuation become more prevalent, so does the need for tractable methods for optimal distributed controller design. In this setting, it is necessary to design control policies that utilize minimal communication between spatially distributed subcontroller units, each with access to a limited subset of system information. This design problem can be formulated as an optimal control problem subject to sparsity constraints that specify the measurement sharing and subcontroller communication structure of the system. It is well-known that such constrained optimization problems are non-convex except in special settings (quadratic invariance [1] or funnel causality [2]).
To overcome the computational complexity that arises from the non-convexity of these constrained optimization problems, various convex restrictions or relaxations have been developed, e.g., [3, 4]. One such convex approximation, the System Level Synthesis (SLS) framework [5], has gained significant recent attention. The SLS framework was originally formulated for finite-dimensional discrete-time systems over a finite time horizon (thus ensuring computational tractability) [5], but has since been adapted to various other settings including infinite time-horizons [6, Sec. 4], infinite spatial dimensions [7], and continuous-time dynamics [8]. Most recently, an SLS-like framework for distributed parameter systems over a continuous spatial domain was developed [9]; this result was restricted to finite-time horizons and discrete-time settings which led to tractable numerical solutions to the problem [9].
The SLS framework is characterized by two key components: a synthesis step and an implementation step. The synthesis step relies on the fact that the cost functional of the design problem may be written as an affine function of well-chosen closed loop system mappings [10]. An “implementation” of the controller is then formed as an interconnection of elements that are built from these closed-loop maps, such that the controller inherits structural properties of the closed-loop mappings. (E.g., imposing that the closed-loops are FIR and sparse ensures that the resulting controller can also be implemented with sparse FIR elements.)
In this work, we demonstrate that an SLS-like procedure can recover the solution to the continuous-time, infinite-time-horizon Linear Quadratic Regulator (LQR) problem for a specific infinite-dimensional linear distributed parameter system: the diffusion equation. Although such optimal control problems have been extensively studied [11, 12], [13, Ch. 7], the formulation we present is novel.
Analogous to the finite-dimensional SLS framework, we demonstrate that (1) the cost functional of interest can be written as an affine function of certain closed-loop mappings, and (2) the controller can be implemented by an interconnection of systems built from these closed-loop mappings. Although our results are presented for the diffusion equation, the approach presented may be modified to apply to other problem settings. This presents a first step toward developing a general SLS-like framework for distributed parameter systems in continuous-time and infinite time-horizon settings, which we believe will complement the computational results and theory developed in the discrete, finite-time horizon setting [9].
This paper is structured as follows. Section II provides notation and mathematical preliminaries. Section III introduces the diffusion dynamics and formulates the $H_{2}$ design problem of interest. Section IV presents our main result: an equivalent formulation of the proposed $H_{2}$ design problem in terms of the systems “closed-loop maps”, analogous to the SLS framework [5]. This problem is solved analytically in Section V, and the structure and stability properties of the resulting controller implementation are explored. The analytical solution is shown to recover the standard infinite-horizon LQR solution in Section VI. Finally, Section VII illustrates that constraints can be imposed on the closed-loop maps while preserving convexity of the underlying $H_{2}$ problem.
II Notation & Mathematical Preliminaries
Let $\mathcal{X},\mathcal{U}$ denote two Hilbert spaces. $\mathcal{L}(\mathcal{X},\mathcal{U})$ denotes the set of linear operators from $\mathcal{X}$ to $\mathcal{U}$, and $\mathcal{L}(\mathcal{X})=\mathcal{L}(\mathcal{X},\mathcal{X}).$ An operator $B\in\mathcal{L}(\mathcal{X},\mathcal{U})$ is bounded if $\sup_{u\in\mathcal{U}}\frac{\|Bu\|_{\mathcal{X}}}{\|u\|_{\mathcal{U}}}<\infty$. $C^{\dagger}\in\mathcal{L}(\mathcal{U},\mathcal{X})$ denotes the adjoint of $C\in\mathcal{L}(\mathcal{X},\mathcal{U})$.
Let $\mathbb{T}$ denote the unit circle, $\mathbb{R}$ the real numbers, $\mathbb{C}$ the complex numbers, and $\mathbb{Z}$ the set of integers. Two Hilbert spaces of interest are the space $L_{n}^{2}(\mathbb{T})$ of $\mathbb{R}^{n}$-valued functions equipped with inner product $\left<f,g\right>_{L_{n}^{2}(\mathbb{T})}=\int_{\theta\in\mathbb{T}}f^{\top}(\theta)g(\theta)d\theta,$ and the space $\ell_{n}^{2}(\mathbb{Z})$ of $\mathbb{R}^{n}$-valued sequences equipped with inner product $\big<\hat{f},\hat{g}\big>_{\ell^{2}_{n}(\mathbb{Z})}=\sum_{\kappa\in\mathbb{Z}}\hat{f}^{\top}(\kappa)\hat{g}(\kappa).$ We neglect the subscript $n$ when clear from context.
Lower-case letters denote spatio-temporal signals over the domain $0,\infty)\times\mathbb{T},$ e.g.,
$\psi(t,\theta),~~t\in[0,\infty),~\theta\in\mathbb{T}.$ (1)
Such signals can also be viewed as Hilbert-space valued signals over time, which we denote with bold-face lower case letters, e.g.,
${\boldsymbol{\psi}}(t):=\psi(t,\cdot).$ (2)
Let $\mathcal{X}$ be a Hilbert-space and $\boldsymbol{\psi}$}} an $\mathcal{X}$-valued signal. Define the signal ${{\boldsymbol{\psi}}}_{\beta}(t)=e^{-\beta t}{{\boldsymbol{\psi}}}(t).$ If ${{\boldsymbol{\psi}}}_{\beta}\in L_{1}([0,\infty),\mathcal{X})$, then $\boldsymbol{\psi}$}} is Laplace transformable with Laplace transform ${{\boldsymbol{\Psi}}}:\{s\in\mathbb{C};~{\rm Re}(s)\geq\beta\}\rightarrow\mathcal{X}$ given by [14, Ch. 7] ${\boldsymbol{\Psi}}(s):=\int_{0}^{\infty}e^{-st}{\boldsymbol{\psi}}(t)dt.$
$\mathcal{F}$ denotes the Fourier transform of elements of $L^{2}(\mathbb{T}):$
$\hat{f}(j\kappa):=(\mathcal{F}f)(j\kappa):=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}f(\theta)e^{-j\kappa\theta}d\theta,$ (3)
which can be applied to the spatial component of spatio-temporal signals, e.g.,
$\hat{\psi}(t,j\kappa)=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}\psi(t,\theta)e^{-j\kappa\theta}d\theta.$ (4)
Recall that $\mathcal{F}$ is an isometric isomorphism to $\ell^{2}(\mathbb{Z})$.
A measurable function $\hat{A}:\mathbb{Z}\rightarrow\mathbb{C}^{m\times n}$, defines a multiplication operator $M_{\hat{A}}\in\mathcal{L}(\ell^{2}_{n}(\mathbb{Z}),\ell^{2}_{m}(\mathbb{Z}))$
$(M_{\hat{A}}\hat{f})(\kappa)=\hat{A}(\kappa)\hat{f}(\kappa).$ (5)
We refer to $\hat{A}_{\kappa}=\hat{A}(\kappa)$ as the symbol of the operator $M_{\hat{A}}$.
Letting ++ denote the addition operator on the group $\mathbb{T},$ for each $\xi\in\mathbb{T}$, define the translation operator $T_{\xi}\in\mathcal{L}(L^{2}_{n}(\mathbb{T})):$
$(T_{\xi}f)(\theta):=f(\theta+\xi).$ (6)
An operator $B\in\mathcal{L}(L^{2}_{m}(\mathbb{T}),L^{2}_{n}(\mathbb{T}))$ is translation invariant if it commutes with every translation operator, i.e. $T_{\xi}B=BT_{\xi}$ for all $\xi\in\mathbb{T}.$
III Problem Set Up
The controller and plant of interest are linear, time- and spatially-invariant distributed parameter systems over the spatial domain $\mathbb{T}$. These systems can be specified in abstract ODE form [14, 13] as
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+(B{\bf u})(t)$ (7)
${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
where $A$ generates a $C_{0}$-semigroup $\{T_{A}(t)\}$ on $\mathcal{D}(A)\subset L_{n}^{2}(\mathbb{T})$ and $B,C,D$ are elements of $\mathcal{L}(L^{2}(\mathbb{T}))$ spaces of compatible dimensions. We denote the system (7) by $\Sigma(A,B,C,D)$; we say that this system is strictly proper if $D\equiv 0$ and that this system is spatially invariant if each of $A,B,C,D$ is a translation invariant operator.
The semigroup $\{T_{A}(t)\}$ is exponentially stable if there exist constants $m,\eta>0$ for which
$\|T_{A}(t)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}\leq me^{-\eta t},~~\text{for all}~t\geq 0.$ (8)
We define the infinite horizon observability Gramian, $L_{O}$, of the pair $(A,C)$ by [14]
$L_{O}=\int_{0}^{\infty}(T_{A}(t))^{\dagger}C^{\dagger}CT_{A}(t)dt,$ (9)
when this integral converges absolutely in operator norm. A sufficient condition for $L_{O}$ to be well-defined is that $C$ is bounded and that $\{T_{A}(t)\}$ is exponentially stable. We leverage $L_{O}$ to define the $H_{2}$ norm of a linear distributed parameter system.
Definition 1
When $B$ is bounded and the infinite horizon observability Gramian $L_{O}$ is trace class, we define the $H_{2}$ norm of the system $\Sigma(A,B,C,D)$ as
$\|\Sigma(A,B,C,D)\|_{H_{2}}=\tfrac{1}{2\pi}{\rm Tr}(B^{\dagger}L_{O}B).$ (10)
Note that boundedness of $L_{O}$ is sufficient for this operator to be trace class; combined with the boundedness of $B$, this ensures the metric (10) will be finite.
We additionally utilize operator-valued transfer function representations of systems [14, Ch.7]. Let $\rho(A)$ denote the resolvent set and $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$ denote the domain of the operator $A$. For $s\in\rho(A)$, the resolvent operator $R(s;A)=(sI-A)^{-1}$ is well-defined on $\mathcal{D}(A)$ by
$R(s;A)\psi_{0}=\int_{0}^{\infty}e^{-st}T_{A}(t)\psi_{0}dt,~\psi_{0}\in\mathcal{D}(A).$ (11)
$\rho(A)$ will contain a set of the form $\{s\ \in\mathbb{C};~{\rm Re}(s)\geq\omega\}$, where $\omega$ defines a growth bound on the semigroup $\{T_{A}(t)\}$. When this semigroup is exponentially stable, one can take $\omega=0$ [14, Ch. 2]. Then the operator-valued transfer function representation of $\Sigma(A,B,C,D)$ is given by
$C(sI-A)^{-1}B+D:\mathcal{D}(A)\rightarrow L^{2}(\mathbb{T}),~~s\in\rho(A).$ (12)
We say this transfer function is strictly proper if $\lim_{s\rightarrow\infty}\left(C(sI-A)^{-1}B+D\right)=0,$ i.e. if $D=0$.
When $\Sigma(A,B,C,D)$ is spatially invariant, each of $A,B,C,D$ can be transformed to a multiplication operator on $\ell^{2}(\mathbb{Z})$ spaces of appropriate dimensions, e.g., $M_{\hat{A}}=\mathcal{F}^{-1}A\mathcal{F}$. In this case, (17) can be represented as
$\tfrac{d}{dt}\hat{{\boldsymbol{\psi}}}(t)$ $=(M_{\hat{A}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{B}}\hat{\bf u})(t)$ (13)
$\hat{\bf z}(t)$ $=(M_{\hat{C}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{D}}\hat{\bf u})(t),$
and its transfer function can also be represented as a multiplication operator [14, Ch.7]:
$\mathcal{F}^{-1}\left(C(sI-A)^{-1}B+D\right)\mathcal{F}=M_{\hat{G}(s)}$ (14)
$\hat{G}_{\kappa}(s)=\hat{C}_{\kappa}(sI-\hat{A}_{\kappa})^{-1}\hat{B}_{\kappa}+\hat{D}_{\kappa}.$
III-A Controller Design for the Diffusion Equation
We focus here on a specific plant - the diffusion equation - noting that our approach could be generalized to other linear distributed parameter systems. The diffusion equation over $\mathbb{T}$ with fully distributed actuation $u$ and disturbance $w$ is described by the partial differential equation (PDE)
$\partial_{t}\psi(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)+u(t,\theta)+w(t,\theta),~t\in[0,\infty),~\theta\in\mathbb{T},$ (15)
where $\alpha>0$ is the constant of diffusivity. The spatio-temporal signal
${z}(t,\theta)=\left[\begin{array}[]{c}\gamma{\psi}(t,\theta)\\ {u}(t,\theta)\end{array}\right],$ (16)
specifies a performance output of interest that provides a uniform penalty across the spatial variable, with $\gamma>0$ determining a tradeoff between cost of state and cost of control. The dynamics (15)-(16) can be written as
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+{\bf u}(t)+{\bf w}(t)$ (17)
${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
where $A\in\mathcal{L}(L_{1}^{2}(\mathbb{T}))$ is defined by
$(A\psi)(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)$ (18)
and generates a $C_{0}$- semigroup $\{T_{A}(t)\}$ on
$\{f\in L^{2}(\mathbb{T});~f^{\prime},f^{\prime\prime}\in L^{2}(\mathbb{\mathbb{T}})\}=:\mathcal{D}(A)\subset L^{2}(\mathbb{T}),$ (19)
with derivatives, $f^{\prime},f^{\prime\prime}$, in (19) interpreted weakly, and $C,D\in\mathcal{L}(L_{1}^{2}(\mathbb{T}),L_{2}^{2}(\mathbb{T}))$ are the bounded operators
$(C\psi)(t,\theta)=\left[\begin{array}[]{c}\gamma\psi(t,\theta)\\ 0\end{array}\right],~(Du)(t,\theta)=\left[\begin{array}[]{c}0\\ u(t,\theta)\end{array}\right].$ (20)
We denote the operator-valued transfer function for (17) by
$P(s)=C(sI-A)^{-1}+D,$ (21)
which is well-defined on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\subset\rho(A)$ since 0 defines a growth bound on the semigroup $\{T_{A}(t)\}.$ $\Sigma(A,I,C,D)$ is a spatially-invariant system that is equivalently characterized by its transformation under $\mathcal{F}$ as
$\tfrac{d}{dt}\hat{\psi}(t,\kappa)=\underbrace{-\alpha\kappa^{2}}_{\hat{A}_{\kappa}}\hat{\psi}(t,\kappa)+\hat{u}(t,\kappa)+\hat{w}(t,\kappa),$ (22)
$\hat{z}(t,\kappa)=\underbrace{\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right]}_{\hat{C}_{\kappa}}\hat{\psi}(t,\kappa)+\underbrace{\left[\begin{array}[]{c}0\\ 1\end{array}\right]}_{\hat{D}_{\kappa}}\hat{u}(t,\kappa),~~t\in[0,\infty),~\kappa\in\mathbb{Z}$
or in transfer function form as $\mathcal{F}^{-1}P(s)\mathcal{F}=M_{\hat{P}(s)},$
$\hat{P}_{\kappa}(s)=\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right](s+\alpha\kappa^{2})^{-1}+\left[\begin{array}[]{c}0\\ 1\end{array}\right].$ (23)
We consider a state feedback control policy defined by a transfer function
${\bf U}(s)=K(s){\boldsymbol{\Psi}}(s),$ (24)
defined on some set $\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$. The feedback interconnection of (17) and (24) is a well-defined transfer function on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\cap\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$ [14, Ex. 7.3], which we denote by $\mathbb{F}(P;K)(s).$
Definition 2
We say that the control policy (24) internally stabilizes the system (17) if $\mathbb{F}(P;K)(s)$ is the operator-valued transfer function of a system that generates an exponentially stable $C_{0}$-semigroup $\{T_{A_{\rm c.l.}}(t)\}$ and $\|\mathbb{F}(P;K)\|_{H_{2}}<\infty$.
We now state formally the $H_{2}$ design problem of interest:
$\inf_{K}~~\|\mathbb{F}(P;K)\|_{H_{2}}$ (25)
$~{\rm s.t.}~~K~\text{spatially-invariant}$
$~~~~~~K~\text{internally stabilizing}.$
As in the finite-dimensional setting, this $H_{2}$ problem will recover the solution to the Linear Quadratic Regulator (LQR) problem for the diffusion equation [15, 14]:
$\inf_{u}~\int_{0}^{\infty}\left<{\boldsymbol{\psi}}(t),\gamma^{2}{\boldsymbol{\psi}}(t)\right>_{L^{2}(\mathbb{T})}+\left<{\bf u}(t),{\bf u}(t)\right>_{L^{2}(\mathbb{T})}dt$ (26)
[…]
IV Closed-Loop Formulation of $H_{2}$ Problem
In this section, we demonstrate that the controller design problem (LABEL:eq:H2_formulation) can be formulated as an optimization problem over certain closed-loop mappings and admits analytical solutions in the spatial frequency domain. We will illustrate that this formulation allows for spatio-temporal constraints to be imposed in a convex manner in Section VII.
We isolate the closed-loop transfer functions from $w$ to $\psi$ and $u$ in the interconnection of (17) and (24) to obtain:
${\boldsymbol{\Psi}}(s)$ $=(I-(sI-A)^{-1}K(s))^{-1}(sI-A)^{-1}{\bf W}(s)$ (27)
$=(sI-A-K(s))^{-1}{\bf W}(s)$
$=:\Phi^{\psi}(s){\bf W}(s)$
${\bf U}(s)$ $=K(s)(I-(sI-A)^{-1}K(s))^{-1}(sI-A)^{-1}{\bf W}(s)$
$=K(s)(sI-A-K(s))^{-1}{\bf W}(s)$
$=:\Phi^{u}(s){\bf W}(s).$
The operator-valued transfer functions $\Phi^{\psi}(s),\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ are analogous to the “system-level maps” introduced for finite dimensional discrete-time systems in [5]. The closed-loop transfer function from $w$ to $z$ can be written in terms of $\Phi^{\psi}(s),\Phi^{u}(s)$ as
${\bf Z}(s)$ $=\mathbb{F}(P;K)(s){\bf W}(s)$ (28)
$=\left(C\Phi^{\psi}(s)+D\Phi^{u}(s)\right){\bf W}(s).$
We parameterize the set of internally stabilizing, time- and spatially-invariant state feedback controllers for (17) in terms of $\Phi^{\psi}(s)$ and $\Phi^{u}(s)$.
Lemma 1
Let $K$ be a spatially-invariant and internally stabilizing controller for (17). Then the operator-valued transfer functions (27) are spatially-invariant, strictly proper, well-defined on $\{s\in\mathbb{C};~{\rm Re}(s)\geq 0\},$ and satisfy
$\sup_{{\rm Re}(s)\geq 0}~\|\Phi^{\psi}(s)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}<\infty,$ (29)
$\sup_{{\rm Re}(s)\geq 0}~\|\Phi^{u}(s)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}<\infty.$
Condition (LABEL:eq:Hinf_condition) can be equivalently verified pointwise in the spatial frequency domain as:
$\sup_{{\rm Re}(s)\geq 0}\sup_{\kappa\in\mathbb{Z}}~\left|\hat{\Phi}^{\psi}_{\kappa}(s)\right|<\infty,~\sup_{{\rm Re}(s)\geq 0}\sup_{\kappa\in\mathbb{Z}}~\left|\hat{\Phi}^{u}_{\kappa}(s)\right|<\infty.$ (30)
Proof:
See Appendix -C. ∎
Given the analogy to the definition for finite dimensional systems, we say that an operator-valued transfer function is an element of $H_{\infty}$ if it satisfies the condition (LABEL:eq:Hinf_condition).
Lemma 2
Let $\Phi^{\psi}(s)$ and $\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ be defined by (27) for some spatially-invariant and internally stabilizing controller $K(s).$ Then, for all $\eta,\vartheta\in L^{2}(\mathbb{T})$ and $s\in\{s\in\mathbb{C};{\rm Re}(s)\geq 0\}$,
$\left<\eta,\left((s+1)I-(A+I)\right)\Phi^{\psi}(s)\vartheta\right>_{L^{2}(\mathbb{T})}$ (31)
$\hskip 72.26999pt-\left<\eta,\Phi^{u}(s)\vartheta\right>_{L^{2}(\mathbb{T})}=\left<\eta,\vartheta\right>_{L^{2}(\mathbb{T})}.$
Equivalently,
$\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,~\forall\kappa\in\mathbb{Z}.$ (32)
Proof:
See Appendix -D. ∎
Lemma 3
Let ${\Phi}^{\psi},{\Phi}^{u}\in H_{\infty}\subset\mathcal{L}(L^{2}(\mathbb{T}))$ be spatially-invariant, strictly proper, and satisfy (32). Select a spatially-invariant transfer function [^1] Existence of such a $G$ will be verified in Section V-A. $G$ for which $G^{-1}(s)\in H_{\infty}$, $\lim_{s\to\infty}s\hat{G}_{\kappa}^{-1}(s)=1~\forall\kappa\in\mathbb{Z}$, and $G^{-1}(s)(sI-A)\in H_{\infty}$. Then,
${\bf v}(s)$ $={{\boldsymbol{\Psi}}}(s)+\left(I-G(s){\Phi}^{\psi}(s)\right){\bf v}(s)$ (33)
${\bf u}(s)$ $=G(s){\Phi}^{u}(s){\bf v}(s)$
defines an internally stabilizing controller for (17). This implementation (33) is well-posed and the internal signal $\bf v$} remains bounded for all bounded signals $\bf n$} and $\bf w$} as depicted in Figure 1. The input-output map (transfer function) corresponding to (33) is
${\bf U}=\mathcal{F}^{-1}M_{\hat{K}(s)}\mathcal{F}{{\boldsymbol{\Psi}}},~~\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}.$ (34)
Proof:
See Appendix -E. ∎
Remark 1
In the finite-dimensional, discrete-time setting, the analogy of the system $G$ in (33) is simply the one-time-step delay operator. More care must be taken to select this system in the infinite-dimensional setting due to the potential unboundedness of the operator $A$. In this case, the closed-loop map from additive noise to the state to the internal control signal $v$ is given by $(sI-A)G^{-1}(s)$ (see (44)), which may also be unbounded unless $G$ is carefully selected.
Figure 1: Block diagram of the dynamic controller implementation (33) with noisy measurements and exogenous disturbance: $w$ is additive noise to the state $\psi$, and $w$ is additive noise to the control signal $u$.
Lemmas 1-3 demonstrate that the set of all internally stabilizing and spatially-invariant controllers for (17) can be characterized by an affine subspace of spatially-invariant transfer functions $\Phi^{\psi},\Phi^{u}.$ With this, we are able to state our main result.
Theorem 1
The following is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation) for the diffusion equation:
$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (35)
$~~~~~{\rm s.t.}~~~~~~\Phi^{\psi},\Phi^{u}\text{ spatially invariant, strictly proper}$
[…]
V Analytical Solution to Closed-Loop Formulation of $H_{2}$ Problem
Spatial invariance properties allow us to solve (LABEL:eq:CL_formulation) in the spatial frequency domain. By the Placherel theorem, formulation (LABEL:eq:CL_formulation) of the $H_{2}$ design problem for the diffusion equation is equivalent to the optimization problem:
$\inf_{{\hat{\Phi}^{\psi}(s)},{\hat{\Phi}^{u}(s)}}~\left\|M_{\hat{C}}M_{\hat{\Phi}^{\psi}(s)}+M_{\hat{D}}M_{\hat{\Phi}^{u}(s)}\right\|_{H_{2}}$ (36)
$~~~~~{\rm s.t.}~~~~~~\hat{\Phi}^{\psi},\hat{\Phi}^{u}\text{ strictly~proper,~spatially~invariant},$
[…]
Note that the cost and constraints of (LABEL:eq:CLOpt_Frequency) “decouple” in the spatial frequency variable $\kappa.$ Thus, the solution to (LABEL:eq:CLOpt_Frequency) can be obtained by solving the family of $H_{2}$ problems, parameterized by $\kappa\in\mathbb{Z}:$
$\inf_{\hat{\Phi}^{\psi}_{\kappa},\hat{\Phi}^{u}_{\kappa}}\left\|\left[\begin{array}[]{cc}\gamma&0\\ 0&1\end{array}\right]\left[\begin{array}[]{c}\hat{\Phi}^{\psi}_{\kappa}\\ \hat{\Phi}^{u}_{\kappa}\end{array}\right]\right\|_{H_{2}}$ (37e)
$~~{\rm s.t.}~~~\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ (37f)
$~~~~~~~~\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,$ (37g)
provided that the family of solutions $\{\hat{\Phi}^{\psi}_{\kappa}\},~\{\hat{\Phi}^{u}_{\kappa}\}$ satisfy the condition (LABEL:eq:Hinf_frequency) which ensures $\Phi^{\psi},\Phi^{u}\in H_{\infty}$.
For each fixed $\kappa\in\mathbb{Z},$ (37) is a standard $H_{2}$ design problem for a finite-dimensional system with an implicit affine subspace constraint (37g). To obtain an analytical solution to this problem, we first replace the implicit constraint (37g) with an analogous explicit constraint.
Lemma 4
Fix $\kappa\in\mathbb{Z}.$ The single-input single-output (SISO) transfer functions $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ satisfy $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ and (37g) if and only if they are of the form
$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+1}\left(1+\hat{\rho}_{\kappa}(s)\right),$ (38)
$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-1}{s+1}+\frac{s+\alpha\kappa^{2}}{s+1}\hat{\rho}_{\kappa}(s)$
for some SISO transfer function $\hat{\rho}_{\kappa}(s)\in RH_{2}.$
Proof:
See Appendix -F. ∎
We leverage the parameterization (38) and properties of the $H_{2}$ norm to write (37) as an unconstrained $H_{2}$ design problem over a single SISO transfer function decision variable $\hat{\rho}_{\kappa}.$
Proposition 1
The optimal solution $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ to the problem (37) can be recovered from the solution $\hat{\rho}_{\kappa}$ to the optimization problem
$\inf_{\hat{\rho}_{\kappa}\in RH_{2}}~\left\|\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{\alpha\kappa^{2}-1}{s+1}\end{array}\right]+\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{s+\alpha\kappa^{2}}{s+1}\end{array}\right]\hat{\rho}_{\kappa}(s)\right\|_{H_{2}}$ (39)
through the relation (38). The optimal solution to (39) is
$\hat{\rho}_{\kappa}^{\rm opt.}(s)=\frac{1-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (40)
and thus the optimal solution to (37) is given by
$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (41)
$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}.$
Proof:
See Appendix -A. ∎
V-A Implementation of Controller via Closed-Loop Solution
We examine the structure of the controller implementation (33) that results from the optimal solution (41). We select $G(s)=\mathcal{F}^{-1}M_{\hat{G}(s)}\mathcal{F}$ where
$\hat{G}_{\kappa}(s)=s+1+\kappa^{2},$ (42)
to satisfy all conditions of Lemma 3. In this case, (33) may be specified in spatial frequency $\kappa$ as
$\hat{v}(s,\kappa)$ $=\hat{\psi}(s,\kappa)+\left(1-\hat{G}_{\kappa}(s)\tfrac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}\right)\hat{v}(s,\kappa)$ (43)
$\hat{u}(s,\kappa)$ $=\hat{G}_{\kappa}(s)\tfrac{\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}\hat{v}(s,\kappa),$
where $\hat{G}$ is defined in (42). We note that the system $\mathcal{F}(I-G(s)\Phi^{\psi}(s))\mathcal{F}^{-1}$ with symbol $\left(1-\hat{G}_{\kappa}(s)\tfrac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}\right)$ is strictly proper, so that the feedback loop in the control implementation in Fig. 1 is well-defined. Using the affine identity in (32) together with routine calculations yields the following closed-loop transfer matrices from $(\hat{n}_{\kappa}(s),\hat{w}_{\kappa}(s))$ to $(\hat{\psi}_{\kappa}(s),\hat{u}_{\kappa}(s),\hat{v}_{\kappa}(s))$ in the controller implementation:
$\begin{bmatrix}\hat{\psi}_{\kappa}\\ \hat{u}_{\kappa}\\ \hat{v}_{\kappa}\end{bmatrix}=\begin{bmatrix}\hat{\Phi}^{u}_{\kappa}(s)&\hat{\Phi}^{\psi}_{\kappa}(s)\\ (s+\alpha\kappa^{2})\hat{\Phi}^{u}_{\kappa}(s)&\hat{\Phi}^{u}_{\kappa}(s)\\ (s+\alpha\kappa^{2})\hat{G}^{-1}_{\kappa}(s)&\hat{G}^{-1}_{\kappa}(s)\end{bmatrix}\begin{bmatrix}\hat{n}_{\kappa}\\ \hat{w}_{\kappa}\\ \end{bmatrix}$, (44)
each of which is verified to be the symbol of a multiplication operator in $H_{\infty}$ for our selected $\hat{G}$. We numerically compute the spatio-temporal impulse responses (Green’s functions) of the blocks in Fig. 1 and plot their dynamic and static components in Fig. 2. We note that the kernels of the dynamic components are “more localized” for smaller time values and appear to spread out spatially as time increases.
Figure 2: Spatio-temporal convolution kernels corresponding to the transfer functions in Fig. 1. Plot (a) shows the yellow block’s static, $h_{\Phi^{u}}(\xi)$, and dynamic $f_{\Phi^{u}}(t,\xi)$ kernels. Plot (b) shows the blue block’s dynamic kernel $f_{\Phi^{\psi}}(t,\xi)$.
VI Equivalence to LQR Control
We demonstrate that the solution to our $H_{2}$ design problem indeed recovers the solution to the standard infinite-horizon LQR problem for the diffusion equation.
Proposition 2
Let $(\mathcal{F}\psi_{0})(\kappa)=1$ for all $\kappa$.  [^2] Note that the “response” (weak solution) of a system to this initial condition is equal to the Green’s Function of the linear PDE describing the system dynamics, and that $\mathcal{F}$ here is the extension of the Fourier transform to the space of distributions. Then the standard LQR problem (LABEL:eq:LQR) is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation).
Proof:
First, we compute the solution to the LQR formulation (LABEL:eq:LQR) using standard approaches [16, 14]. We solve the operator-valued Ricatti equation
$\left<A\eta,F\vartheta\right>_{L^{2}(\mathbb{T})}+\left<F\eta,A\vartheta\right>_{L^{2}(\mathbb{T})}+\gamma^{2}\left<\eta,\vartheta\right>_{L^{2}(\mathbb{T})}$ (45)
$=\left<F\eta,F\vartheta\right>_{L^{2}(\mathbb{T})}.$
The solution $F$ to (45) is known to be spatially-invariant [16], and the symbol of $\mathcal{F}F\mathcal{F}^{-1}=M_{\hat{F}}$ can be obtained by solving the family of scalar-valued algebraic Riccati equations (AREs) parameterized by $\kappa\in\mathbb{Z}:$
$\alpha\kappa^{2}\hat{F}_{\kappa}-\hat{F}_{\kappa}\alpha\kappa^{2}-\hat{F}_{\kappa}^{2}+\gamma^{2}=0.$ (46)
The unique solution $\hat{F}_{\kappa}>0$ to (46) is $\hat{F}_{\kappa}=-\alpha\kappa^{2}+\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}},$ and the optimal state feedback gain is thus given by
$K=\mathcal{F}M_{\hat{K}}\mathcal{F}^{-1},~~\hat{K}_{\kappa}=-\hat{F}_{\kappa}=\alpha\kappa^{2}-\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}}.$ (47)
We next obtain the input-output mapping specified by the controller implementation (43) using (34) as
$\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}=\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}.$ (48)
This is indeed equal to the static state feedback gain $\hat{K}$ obtained from the Riccati equation solution. Details on the equivalence of the cost of the two formulations (LABEL:eq:CL_formulation) and (LABEL:eq:LQR) are provided in Appendix -B. ∎
We know that $K$ will take the form of a spatial convolution operator
$u(t,\theta)=\int_{\xi\in\mathbb{T}}k(\theta-\xi)\psi(t,\xi)dt,$ (49)
where $k(\theta)=(\mathcal{F}^{-1}\hat{K})(\theta).$ Numerical integration shows that $k(\theta)$ exhibits spatial decay away from the origin (see Fig. 2).
We compare the static controller $K$ to the dynamic controller implementation (43) depicted in Fig. 1. As previously noted, the transfer function $\left(I-G(s){\Phi}^{\psi}(s)\right)$ is strictly proper, thus its corresponding convolution kernel has only a dynamic part, which we denote as $f_{\Phi^{\psi}}(t,\theta)$. The transfer function $G(s)\Phi^{u}(s)$ has a corresponding convolution kernel with both a static and dynamic part, which we denote as $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta)$, respectively. The static (feedthrough) component of $G(s)\Phi^{u}(s)$ is obtained by evaluating its Fourier symbol in the limit as $s\to\infty$,
$\mathcal{F}^{-1}\lim_{s\to\infty}\hat{G}(s,\kappa)\hat{\Phi}^{u}(s,\kappa)=h_{\Phi^{u}}(\theta)=(\mathcal{F}^{-1}\hat{K})(\theta),$
which shows that the static part of the implementation (43) recovers the static LQR gain, as expected. The “dynamic components” of the systems $G(s)\Phi^{u}(s)$ and $\left(I-G(s){\Phi}^{\psi}(s)\right)$ cancel each other out, leaving only the static feedback policy $K.$ This is confirmed in the plot of $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta),f_{\Phi^{\psi}}(t,\theta)$ shown in Fig. 2.
VII Incorporation of Spatio-Temporal Decay Rate Constraints
Without constraints, the closed-loop formulation of the $H_{2}$ problem (LABEL:eq:CL_formulation) is equivalent to a standard LQR problem. The advantage of this closed-loop formulation is that it remains convex when spatio-temporal constraints are incorporated. A similar problem has been studied in [2], where necessary and sufficient conditions for convexity were dependent on “funnel causality” properties of the underlying system and desired constraint. Notably, our formulation does not rely on these conditions and thus admits a broader range of constraints while preserving convexity.
An added constraint that $I-G\Phi^{\psi},G\Phi^{u}\in\mathcal{S}$ enforces a structural constraint on the components of our controller. In this case, formulation (LABEL:eq:CL_formulation) becomes:
$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (50)
$~~~~~{\rm s.t.}~~~~~~\Phi^{\psi},\Phi^{u}\text{ spatially invariant}$
[…]
$~~~~~~~~~~~~~~~~I-G\Phi^{\psi},G\Phi^{u}\in\mathcal{S}.$
So long as the set $\mathcal{S}$ defines a convex constraint on $\Phi^{u},\Phi^{\psi}$ the constrained problem (LABEL:eq:constrained) will be convex.
Of particular interest are constraints that specify a degree of spatio-temporal locality on $\Phi^{\psi}$ and $\Phi^{u}$. In the finite-dimensional, discrete-time, finite-horizon setting, such constraints take the form of spatio-temporal FIR constraints on the closed-loop system [5]. In the continuous spatial and temporal domain setting, we believe it is more natural to impose a level of decay, or “approximate locality”. This choice might be viewed as analogue to the “leaky” SLS problem presented in [17].
One possible spatio-temporal constraint is when $\mathcal{S}$ is the set of linear PDEs whose Green’s functions $\rho(\cdot,\cdot)$ satisfy
$\|\rho(\cdot,t)-\rho_{T}(\cdot,t)\|_{L_{2}(\mathbb{T})}<\epsilon(t)$
where $\rho_{T}$ is the spatial truncation of $\rho$ given by
$\rho_{T}(\theta,\cdot)=\begin{cases}\rho(\theta,\cdot)&|\theta|<r\\ 0&{\rm otherwise}\end{cases}$
and $r\leq\pi$ is the truncation radius that encodes communication limitations. This constraint ensures that the ‘tails’ of $\rho$ are sufficiently small, potentially enforcing spatial locality. Choosing $\epsilon(t)$ to grow in $t$ allows $\rho$ to be less localized in space as time increases. This set $\mathcal{S}$ is convex, and guarantees that a spatially truncated version of controller will perform well using small gain arguments.
Obtaining analytical and numerical solutions to constrained problems of the form (LABEL:eq:constrained) is the subject of ongoing work. While analytical solutions will require clever formulations of the constraints, numerical solutions may leverage approaches such as those presented in [9] and benefit from computational approaches like [18].
VIII Conclusion
It was shown that the infinite-horizon LQR problem for the diffusion equation over the unit circle can be reformulated as an $H_{2}$ problem over the system’s closed-loop responses, analogous to the finite-dimensional System Level Synthesis methodology. Implementation of the resulting controller using the closed-loop responses was shown to preserve stability of the closed-loop system. It was further illustrated that this novel formulation admits constraints on the closed-loop responses while preserving convexity. Generalizations of the presented results to a class of linear distributed parameter systems over both compact and unbounded spatial domains are the subject of ongoing work. Of particular interest is comparing analytical results from this framework to the computational approach of [9].

### v2

@@
 Abstract
-The continuous-time, infinite horizon LQR problem for the diffusion equation over the unit circle with fully distributed actuation is considered.
-It is well-known that the solution to this problem can be obtained from the solution to an operator-valued algebraic Riccati equation.
-Here, it is demonstrated that this solution can be equivalently obtained by solving an $H_{2}$ control problem through a closed-loop design procedure that is analogous to the “System Level Synthesis” methodology previously developed for systems over a discrete spatial domain and/or over a finite time horizon.
-The presented extension to the continuous spatial domain and continuous and infinite-horizon time setting admits analytical solutions that may complement computational approaches for discrete or finite-horizon settings.
-It is further illustrated that spatio-temporal constraints on the closed-loop responses can be incorporated into this new formulation in a convex manner.
+The optimal controller design problem for a linear, first-order spatially-invariant distributed parameter system is considered.
+Through a case study of the Linear Quadratic Regulator (LQR) problem for the diffusion equation over the torus, it is illustrated that the optimal controller design problem can be equivalently formulated as an optimization problem over the system’s closed-loop mappings, analogous to the System Level Synthesis framework.
+This reformulation is solved analytically to recover the LQR for the diffusion equation, and an internally stable implementation of this controller is recovered from the optimal closed-loop mappings.
+It is further demonstrated that a class of spatio-temporal constraints on the closed-loop maps can be imposed on this closed-loop formulation while preserving convexity.
 I Introduction
-As large-scale spatially-distributed systems with immense amounts of sensing and actuation become more prevalent, so does the need for tractable methods for optimal distributed controller design.
-In this setting, it is necessary to design control policies that utilize minimal communication between spatially distributed subcontroller units, each with access to a limited subset of system information.
-This design problem can be formulated as an optimal control problem subject to sparsity constraints that specify the measurement sharing and subcontroller communication structure of the system.
-It is well-known that such constrained optimization problems are non-convex except in special settings (quadratic invariance [1] or funnel causality [2]).
-To overcome the computational complexity that arises from the non-convexity of these constrained optimization problems, various convex restrictions or relaxations have been developed, e.g., [3, 4].
-One such convex approximation, the System Level Synthesis (SLS) framework [5], has gained significant recent attention.
-The SLS framework was originally formulated for finite-dimensional discrete-time systems over a finite time horizon (thus ensuring computational tractability) [5], but has since been adapted to various other settings including infinite time-horizons [6, Sec.
-4], infinite spatial dimensions [7], and continuous-time dynamics [8].
-Most recently, an SLS-like framework for distributed parameter systems over a continuous spatial domain was developed [9]; this result was restricted to finite-time horizons and discrete-time settings which led to tractable numerical solutions to the problem [9].
-The SLS framework is characterized by two key components: a synthesis step and an implementation step.
-The synthesis step relies on the fact that the cost functional of the design problem may be written as an affine function of well-chosen closed loop system mappings [10].
-An “implementation” of the controller is then formed as an interconnection of elements that are built from these closed-loop maps, such that the controller inherits structural properties of the closed-loop mappings.
-(E.g., imposing that the closed-loops are FIR and sparse ensures that the resulting controller can also be implemented with sparse FIR elements.)
-In this work, we demonstrate that an SLS-like procedure can recover the solution to the continuous-time, infinite-time-horizon Linear Quadratic Regulator (LQR) problem for a specific infinite-dimensional linear distributed parameter system: the diffusion equation.
-Although such optimal control problems have been extensively studied [11, 12], [13, Ch.
-7], the formulation we present is novel.
-Analogous to the finite-dimensional SLS framework, we demonstrate that (1) the cost functional of interest can be written as an affine function of certain closed-loop mappings, and (2) the controller can be implemented by an interconnection of systems built from these closed-loop mappings.
-Although our results are presented for the diffusion equation, the approach presented may be modified to apply to other problem settings.
-This presents a first step toward developing a general SLS-like framework for distributed parameter systems in continuous-time and infinite time-horizon settings, which we believe will complement the computational results and theory developed in the discrete, finite-time horizon setting [9].
-This paper is structured as follows.
-Section II provides notation and mathematical preliminaries.
-Section III introduces the diffusion dynamics and formulates the $H_{2}$ design problem of interest.
-Section IV presents our main result: an equivalent formulation of the proposed $H_{2}$ design problem in terms of the systems “closed-loop maps”, analogous to the SLS framework [5].
-This problem is solved analytically in Section V, and the structure and stability properties of the resulting controller implementation are explored.
-The analytical solution is shown to recover the standard infinite-horizon LQR solution in Section VI.
-Finally, Section VII illustrates that constraints can be imposed on the closed-loop maps while preserving convexity of the underlying $H_{2}$ problem.
+As large-scale spatially-distributed systems with immense sensing and actuation become more prevalent, so does the need for tractable methods for optimal distributed controller design.
+One motivating example is the development of RF metamaterials for quantum computing, beamforming, and secure communications [23, 13, 21], which are modeled as continnum structures governed by PDEs.
+To leverage distributed sensing and actuation in these and other PDE systems, it is necessary to design control policies that use minimal communication between distributed subcontroller units, each with access to a limited subset of system information.
+This requires solving an optimal control problem subject to controller locality constraints, which is known to be non-convex except in special settings [18, 3].
+To overcome the computational complexity of such constrained optimization problems, various convex restrictions or relaxations have been developed, e.g., [19, 24].
+Of particular significance is the System Level Synthesis (SLS) framework [22], which is characterized by (1) a synthesis step to formulate the design problem and constraints as affine in the system’s closed-loop mappings; and (2) an implementation step to construct an internally stable implementation of the controller from an interconnection of the optimal closed-loop maps.
+This implementation ensures that the controller inherits any structural properties constrained on the closed-loop maps.
+For example, imposing that the closed-loops are spatially sparse ensures that the resulting controller can also be implemented sparse elements.
+The SLS framework was first formulated for finite-dimensional discrete-time systems over a finite time horizon [22], but has since been adapted to various other settings including infinite time-horizons [1, Sec.
+4], countably infinite spatial dimensions [11], and continuous-time dynamics [12].
+However, these approaches could not be directly applied to PDE systems formulated over a continuous spatial domain, as noted by the recent work of Conger et.
+al.
+[6].
+Indeed, it is well-known that discretized approaches to design continuum controllers may lead to problematic “spillover effects” when the solutions to discretized formulations do not appropriately converge to the true continuum solution [16, 10] leading to excitement of problematic modes or even instability.
+This motivates the need for a development of an SLS framework in such continuous spatial domain settings.
+The work of [6] presents a significant step in this direction.
+Notably, these results were restricted to finite-time horizons and discrete-time settings to enable tractable numerical solutions [6], but it remains unclear whether an analogous implementation of the controller can be obtained without such discrete-time and finite-time-horizon assumptions, while preserving internal stability.
+This work builds on the foundation in [6] to address this gap via an informative case study for control of the diffusion equation.
+The main results presented in this paper are as follows:
+  • The infinite time horizon LQR problem for the diffusion equation over the torus is equivalently formulated as an optimization over the closed-loop mappings, analogous to the finite-dimensional SLS framework (Theorem 1).
+  • This closed-loop formulation is solved analytically to recover the known LQR (Sections IV-C, VI).
+  • An implementation of the optimal controller is constructed from the optimal closed-loop mappings, and is rigorously shown to preserve internal stability of the full infinite-dimensional system (Lemma 3).
+  • A class of spatio-temporal constraints on the closed-loop mappings are shown to preserve convexity of the closed-loop design problem (Section VII).
+This work provides a step toward developing a general SLS-like framework for optimal controller design for distributed parameter systems in continuous-time and infinite time-horizon settings, which will complement the computational results and theory developed in [6].
+Although the diffusion equation itself is not particularly interesting, it presents a proof of concept that can inform the use of similar methodology for other first-order in time linear, constant coefficient PDE systems.
 II Notation & Mathematical Preliminaries
-Let $\mathcal{X},\mathcal{U}$ denote two Hilbert spaces.
-$\mathcal{L}(\mathcal{X},\mathcal{U})$ denotes the set of linear operators from $\mathcal{X}$ to $\mathcal{U}$, and $\mathcal{L}(\mathcal{X})=\mathcal{L}(\mathcal{X},\mathcal{X}).$ An operator $B\in\mathcal{L}(\mathcal{X},\mathcal{U})$ is bounded if $\sup_{u\in\mathcal{U}}\frac{\|Bu\|_{\mathcal{X}}}{\|u\|_{\mathcal{U}}}<\infty$.
+Let $\mathbb{T}$ denote the unit circle, $\mathbb{R}$ the real numbers, $\mathbb{C}$ the complex numbers, and $\mathbb{Z}$ the integers.
+For $\mathcal{X},\mathcal{U}$ two Hilbert spaces, $\mathcal{L}(\mathcal{X},\mathcal{U})$ denotes the set of linear operators from $\mathcal{X}$ to $\mathcal{U}$, and $\mathcal{L}(\mathcal{X})=\mathcal{L}(\mathcal{X},\mathcal{X}).$ An operator $B\in\mathcal{L}(\mathcal{X},\mathcal{U})$ is bounded if $\sup_{u\in\mathcal{U}}\frac{\|Bu\|_{\mathcal{X}}}{\|u\|_{\mathcal{U}}}<\infty$.
 $C^{\dagger}\in\mathcal{L}(\mathcal{U},\mathcal{X})$ denotes the adjoint of $C\in\mathcal{L}(\mathcal{X},\mathcal{U})$.
-Let $\mathbb{T}$ denote the unit circle, $\mathbb{R}$ the real numbers, $\mathbb{C}$ the complex numbers, and $\mathbb{Z}$ the set of integers.
-Two Hilbert spaces of interest are the space $L_{n}^{2}(\mathbb{T})$ of $\mathbb{R}^{n}$-valued functions equipped with inner product $\left<f,g\right>_{L_{n}^{2}(\mathbb{T})}=\int_{\theta\in\mathbb{T}}f^{\top}(\theta)g(\theta)d\theta,$ and the space $\ell_{n}^{2}(\mathbb{Z})$ of $\mathbb{R}^{n}$-valued sequences equipped with inner product $\big<\hat{f},\hat{g}\big>_{\ell^{2}_{n}(\mathbb{Z})}=\sum_{\kappa\in\mathbb{Z}}\hat{f}^{\top}(\kappa)\hat{g}(\kappa).$ We neglect the subscript $n$ when clear from context.
-Lower-case letters denote spatio-temporal signals over the domain $0,\infty)\times\mathbb{T},$ e.g.,
+Two Hilbert spaces of interest are the space $L^{2}(\mathbb{T})$ equipped with inner product $\left<f,g\right>_{L^{2}(\mathbb{T})}=\int_{\theta\in\mathbb{T}}f^{*}(\theta)g(\theta)d\theta,$ and the space $\ell^{2}(\mathbb{Z})$ of sequences equipped with inner product $\big<\hat{f},\hat{g}\big>_{\ell^{2}(\mathbb{Z})}=\sum_{\kappa\in\mathbb{Z}}\hat{f}^{*}(\kappa)\hat{g}(\kappa).$
+We consider spatio-temporal signals over the domain $0,\infty)\times\mathbb{T},$ which we denote by lower-case letters e.g.,
 $\psi(t,\theta),~~t\in[0,\infty),~\theta\in\mathbb{T}.$ (1)
-Such signals can also be viewed as Hilbert-space valued signals over time, which we denote with bold-face lower case letters, e.g.,
-${\boldsymbol{\psi}}(t):=\psi(t,\cdot).$ (2)
-Let $\mathcal{X}$ be a Hilbert-space and $\boldsymbol{\psi}$}} an $\mathcal{X}$-valued signal.
-Define the signal ${{\boldsymbol{\psi}}}_{\beta}(t)=e^{-\beta t}{{\boldsymbol{\psi}}}(t).$ If ${{\boldsymbol{\psi}}}_{\beta}\in L_{1}([0,\infty),\mathcal{X})$, then $\boldsymbol{\psi}$}} is Laplace transformable with Laplace transform ${{\boldsymbol{\Psi}}}:\{s\in\mathbb{C};~{\rm Re}(s)\geq\beta\}\rightarrow\mathcal{X}$ given by [14, Ch.
-7] ${\boldsymbol{\Psi}}(s):=\int_{0}^{\infty}e^{-st}{\boldsymbol{\psi}}(t)dt.$
-$\mathcal{F}$ denotes the Fourier transform of elements of $L^{2}(\mathbb{T}):$
+We assume that such spatio-temporal signals $\psi(t,\cdot)\in L^{2}(\mathbb{T})$ for each $t$ and use bold-face lower case letters to denote the corresponding $L^{2}(\mathbb{T})$ valued signals, e.g., ${\boldsymbol{\psi}}(t):=\psi(t,\cdot).$
+The Fourier transform, $\mathcal{F}$, of elements of $L^{2}(\mathbb{T}):$
 $\hat{f}(j\kappa):=(\mathcal{F}f)(j\kappa):=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}f(\theta)e^{-j\kappa\theta}d\theta,$ (2)
-which can be applied to the spatial component of spatio-temporal signals, e.g.,
-$\hat{\psi}(t,j\kappa)=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}\psi(t,\theta)e^{-j\kappa\theta}d\theta.$ (4)
-Recall that $\mathcal{F}$ is an isometric isomorphism to $\ell^{2}(\mathbb{Z})$.
-A measurable function $\hat{A}:\mathbb{Z}\rightarrow\mathbb{C}^{m\times n}$, defines a multiplication operator $M_{\hat{A}}\in\mathcal{L}(\ell^{2}_{n}(\mathbb{Z}),\ell^{2}_{m}(\mathbb{Z}))$
-$(M_{\hat{A}}\hat{f})(\kappa)=\hat{A}(\kappa)\hat{f}(\kappa).$ (5)
-We refer to $\hat{A}_{\kappa}=\hat{A}(\kappa)$ as the symbol of the operator $M_{\hat{A}}$.
-Letting ++ denote the addition operator on the group $\mathbb{T},$ for each $\xi\in\mathbb{T}$, define the translation operator $T_{\xi}\in\mathcal{L}(L^{2}_{n}(\mathbb{T})):$
-$(T_{\xi}f)(\theta):=f(\theta+\xi).$ (6)
-An operator $B\in\mathcal{L}(L^{2}_{m}(\mathbb{T}),L^{2}_{n}(\mathbb{T}))$ is translation invariant if it commutes with every translation operator, i.e. $T_{\xi}B=BT_{\xi}$ for all $\xi\in\mathbb{T}.$
+can be applied to the spatial component of the spatio-temporal signal $\psi(\cdot,\cdot)$:
+$\hat{\psi}(t,j\kappa)=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}\psi(t,\theta)e^{-j\kappa\theta}d\theta,$ (3)
+such that for each $t,$ $\boldsymbol{\psi}(t,\cdot)\in\ell^{2}(\mathbb{Z}).$
+Letting ++ denote the addition operator on the group $\mathbb{T},$ for each $\xi\in\mathbb{T}$, define the translation operator $T_{\xi}\in\mathcal{L}(L^{2}(\mathbb{T})):$ $(T_{\xi}f)(\theta):=f(\theta+\xi).$ An operator $B\in\mathcal{L}(L^{2}(\mathbb{T}))$ is translation invariant if it commutes with every translation operator, i.e.
+$T_{\xi}B=BT_{\xi}$ for all $\xi\in\mathbb{T}.$ Such an operator can be represented in the Fourier domain as a multiplication operator, $B=\mathcal{F}^{-1}M_{\hat{B}}\mathcal{F}$ with $M_{\hat{B}}\in\mathcal{L}(\ell^{2}(\mathbb{Z}))$ such that
+$(M_{\hat{B}}\hat{f})(\kappa)=\hat{B}(\kappa)\hat{f}(\kappa).$ (4)
+We refer to the measurable function $\hat{B}_{\kappa}=\hat{B}(\kappa)$ as the symbol of the operator $M_{\hat{B}}$.
 III Problem Set Up
@@
 These systems can be specified in abstract ODE form [7, 16] as
-$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+(B{\bf u})(t)$ (7)
-${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
-where $A$ generates a $C_{0}$-semigroup $\{T_{A}(t)\}$ on $\mathcal{D}(A)\subset L_{n}^{2}(\mathbb{T})$ and $B,C,D$ are elements of $\mathcal{L}(L^{2}(\mathbb{T}))$ spaces of compatible dimensions.
-We denote the system (7) by $\Sigma(A,B,C,D)$; we say that this system is strictly proper if $D\equiv 0$ and that this system is spatially invariant if each of $A,B,C,D$ is a translation invariant operator.
-The semigroup $\{T_{A}(t)\}$ is exponentially stable if there exist constants $m,\eta>0$ for which
-$\|T_{A}(t)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}\leq me^{-\eta t},~~\text{for all}~t\geq 0.$ (8)
-We define the infinite horizon observability Gramian, $L_{O}$, of the pair $(A,C)$ by [14]
-$L_{O}=\int_{0}^{\infty}(T_{A}(t))^{\dagger}C^{\dagger}CT_{A}(t)dt,$ (9)
+$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=A{\boldsymbol{\psi}}(t)+B({\boldsymbol{u}}+{\boldsymbol{w}})(t)$ (5)
+${\boldsymbol{z}}(t)$ $=C{\boldsymbol{\psi}}(t)+D{\boldsymbol{u}}(t),$
+where $A,B,C,D$ are spatially-invariant operators, $A$ generates a $C_{0}$-semigroup $\{T_{A}(t)\}$ on $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$, $B\in\mathcal{L}(L^{2}(\mathbb{T}))$, and $C,D\in\mathcal{L}(L^{2}(\mathbb{T}),L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T}))$.
+$\boldsymbol{u,w,\psi}$ and $\boldsymbol{z}$} denote the control signal, external disturbance, state, and performance output signal, respectively.
+We denote the system (5) by $\Sigma(A,B,C,D)$.
+A linear, spatially-invariant state feedback controller for this system takes the form
+${\boldsymbol{u}}(t)=K\boldsymbol{\psi}(t),$ (6)
+where $K\in\mathcal{L}(L^{2}(\mathbb{T}))$ is a translation invariant operator.
+We denote the system resulting from interconnection of (5) with controller $K$ as $\mathbb{F}(P;K)$; this system is also spatially invariant and is described by
+$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A+BK){\boldsymbol{\psi}}(t)+B{\bf w}(t)$ (7)
+${\bf z}(t)$ $=(C+DK){\boldsymbol{\psi}}(t).$
+The $C_{0}$-semigroup $\{T_{A_{\rm c.l.}}(t)\}$ generated by $A_{\rm c.l.}:=A+BK$ is exponentially stable if for some constants $m,\eta>0$,
+$\|T_{A_{\rm c.l.}}(t)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}\leq me^{-\eta t},~~\text{for all}~t\geq 0.$ (8)
+Definition 1
+The infinite horizon observability Gramian, $L_{O}$, of the pair $(A_{\rm c.l.},C)$ is [7]
+$L_{O}=\int_{0}^{\infty}(T_{A_{\rm c.l.}}(t))^{\dagger}C^{\dagger}CT_{A_{\rm c.l.}}(t)dt,$ (9)
 when this integral converges absolutely in operator norm.
-A sufficient condition for $L_{O}$ to be well-defined is that $C$ is bounded and that $\{T_{A}(t)\}$ is exponentially stable.
-We leverage $L_{O}$ to define the $H_{2}$ norm of a linear distributed parameter system.
-Definition 1
-When $B$ is bounded and the infinite horizon observability Gramian $L_{O}$ is trace class, we define the $H_{2}$ norm of the system $\Sigma(A,B,C,D)$ as
-$\|\Sigma(A,B,C,D)\|_{H_{2}}=\tfrac{1}{2\pi}{\rm Tr}(B^{\dagger}L_{O}B).$ (10)
-Note that boundedness of $L_{O}$ is sufficient for this operator to be trace class; combined with the boundedness of $B$, this ensures the metric (10) will be finite.
-We additionally utilize operator-valued transfer function representations of systems [14, Ch.7].
+A sufficient condition for $L_{O}$ to be well-defined is that $C$ is bounded and that $\{T_{A_{\rm c.l.}}(t)\}$ is exponentially stable.
+When $B$ is bounded and $L_{O}$ is trace class, the $\mathcal{H}_{2}$ norm [^1] Since we assumed ${\boldsymbol{w}}(t)\in L^{2}(\mathbb{T})$, we can equivalently interpret $\|\cdot\|_{\mathcal{H}_{2}}$ as the $L^{2}$ norm of the impulse response as outlined in [16, Ch.
+5.1].
+of the system $\mathbb{F}(P;K)=\Sigma(A_{\rm c.l.},B,(C+DK),0)$ is defined by
+$\|\Sigma(A_{\rm c.l.},B,(C+DK),0)\|_{\mathcal{H}_{2}}=\tfrac{1}{2\pi}{\rm Tr}(B^{\dagger}L_{O}B).$ (10)
+When $A_{\rm c.l.}$ generates an exponentially stable $C_{0}$-semigroup and $\|\Sigma(A_{\rm c.l.},B,(C+DK),0)\|_{\mathcal{H}_{2}}<\infty$, we say that the controller $K$ is internally stabilizing.
+III-A Diffusion Dynamics
+We focus on a specific plant - the diffusion equation over $\mathbb{T}$ with fully distributed actuation $u$ and disturbance $w$, described by the partial differential equation (PDE)
+$\partial_{t}\psi(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)+u(t,\theta)+w(t,\theta),~t\in[0,\infty),~\theta\in\mathbb{T},$ (11)
+where $u(t,\cdot),w(t,\cdot),\psi(t,\cdot)\in L^{2}(\mathbb{T})$ and $\alpha>0$ is the constant of diffusivity. The spatio-temporal signal
+${z}(t,\theta)=\left[\begin{array}[]{c}\gamma{\psi}(t,\theta)\\ {u}(t,\theta)\end{array}\right],~z(t,\cdot)\in L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T})$ (12)
+is a performance output that provides a uniform penalty across the spatial domain with $\gamma>0$ determining a tradeoff between cost of state and cost of control.
+The interest in this system is not for its particular physical significance, but rather to serve as a case study for the proposed framework.
+Restricting to the torus is equivalent to enforcing periodic boundary conditions and will admit analytic solutions via Fourier analysis approaches [2].
+The dynamics (11)-(12) can be written in the form (5) as
+$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=A{\boldsymbol{\psi}}(t)+{{\boldsymbol{u}}}(t)+{{\boldsymbol{w}}}(t)$ (13)
+${{\boldsymbol{z}}}(t)$ $=C{\boldsymbol{\psi}}(t)+D{{\boldsymbol{u}}}(t),$
+where $A\in\mathcal{L}(L^{2}(\mathbb{T}))$ is defined by $(A\psi)(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)$ and generates a $C_{0}$- semigroup $\{T_{A}(t)\}$ on
+$\{f\in L^{2}(\mathbb{T});~f^{\prime},f^{\prime\prime}\in L^{2}(\mathbb{\mathbb{T}})\}=:\mathcal{D}(A)\subset L^{2}(\mathbb{T}),$ (14)
+with derivatives, $f^{\prime},f^{\prime\prime}$, in (14) interpreted weakly, and $C,D\in\mathcal{L}(L^{2}(\mathbb{T}),L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T}))$ are the bounded operators
+$(C\psi)(t,\theta)=\left[\begin{array}[]{c}\gamma\psi(t,\theta)\\ 0\end{array}\right],~(Du)(t,\theta)=\left[\begin{array}[]{c}0\\ u(t,\theta)\end{array}\right].$
+$\Sigma(A,I,C,D)$ is a spatially-invariant system, equivalently characterized by its transformation under $\mathcal{F}$ as
+$\tfrac{d}{dt}\hat{\psi}(t,\kappa)=-\alpha\kappa^{2}\hat{\psi}(t,\kappa)+\hat{u}(t,\kappa)+\hat{w}(t,\kappa),$ (15)
+$\hat{z}(t,\kappa)=\underbrace{\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right]}_{\hat{C}_{\kappa}}\hat{\psi}(t,\kappa)+\underbrace{\left[\begin{array}[]{c}0\\ 1\end{array}\right]}_{\hat{D}_{\kappa}}\hat{u}(t,\kappa),~t\in[0,\infty),~\kappa\in\mathbb{Z}.$
+Note that for each fixed $\kappa\in\mathbb{Z},$ (LABEL:eq:open_loop_kappa) is a system with finitely many inputs and outputs.
+III-B Optimal Control Formulations
+We consider the Linear Quadratic Regulator (LQR) problem for the diffusion equation (13) [4, 7]:
+$\inf_{u}~\int_{0}^{\infty}\left<{\boldsymbol{\psi}}(t),\gamma^{2}{\boldsymbol{\psi}}(t)\right>_{L^{2}(\mathbb{T})}+\left<{{\boldsymbol{u}}}(t),{{\boldsymbol{u}}}(t)\right>_{L^{2}(\mathbb{T})}dt$ (16)
+$~{\rm s.t.}~~\partial_{t}\psi(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)+u(t,\theta),~{{\boldsymbol{\psi}}}(0)=\psi_{0}.$
+The solution $K^{\rm opt}$ to (LABEL:eq:LQR) is independent of initial condition $\psi_{0}\in L^{2}(\mathbb{T})$ and it will be shown in Section VI that $K^{\rm opt}$ is the solution to the following $\mathcal{H}_{2}$ problem:
+$K^{\rm opt}=$ $\operatorname*{arg\,min}_{K}~~\|\mathbb{F}(P;K)\|_{\mathcal{H}_{2}}$ (17)
+$~~~~{\rm s.t.}~~~~K~\text{spatially-invariant}$
+$~~~~~~~~~~~~K~\text{internally stabilizing}.$
+Remark 1
+In some cases (LABEL:eq:LQR) and (17) result in the same $K^{\rm opt}$, but this does not always hold for PDE systems.
+E.g., if operation by $K$ requires convolution with a delta distribution, the impulse response of the closed-loop system will not be $L^{2}(\mathbb{T})$; thus, the $\mathcal{H}_{2}$ problem is not well-posed, although the LQR problem is well-posed with finite cost $\forall~\psi_{0}\in\mathcal{D}(A)$.
+IV Closed-Loop Formulation of $\mathcal{H}_{2}$ Problem
+In this section, we demonstrate that the controller design problem (17) can be formulated as an optimization problem over certain closed-loop mappings and admits analytical solutions in the spatial frequency domain.
+We will illustrate that this formulation allows for spatio-temporal constraints to be imposed in a convex manner in Section VII.
+IV-A Transfer Functions
+To introduce closed-loop mappings, we use operator-valued transfer function representations of systems [7, Ch.7] to denote mappings between temporally Laplace transformed variables.
+Under certain technical conditions [7, Ch.
+7], for some positive constant $\beta$, the Laplace transform of a signal $\boldsymbol{\psi}$}}, ${{\boldsymbol{\Psi}}}:\{s\in\mathbb{C};~{\rm Re}(s)\geq\beta\}\rightarrow L^{2}(\mathbb{T})$, is well-defined by ${\boldsymbol{\Psi}}(s):=\int_{0}^{\infty}e^{-st}{\boldsymbol{\psi}}(t)dt.$ First let $\Sigma(A,B,C,D)$ be a general system (5).
 Let $\rho(A)$ denote the resolvent set and $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$ denote the domain of the operator $A$.
-For $s\in\rho(A)$, the resolvent operator $R(s;A)=(sI-A)^{-1}$ is well-defined on $\mathcal{D}(A)$ by
-$R(s;A)\psi_{0}=\int_{0}^{\infty}e^{-st}T_{A}(t)\psi_{0}dt,~\psi_{0}\in\mathcal{D}(A).$ (11)
+For $s\in\rho(A)$, the resolvent operator $(sI-A)^{-1}$ is well-defined on $\mathcal{D}(A)$ by
+$(sI-A)^{-1}\psi_{0}:=\int_{0}^{\infty}e^{-st}T_{A}(t)\psi_{0}dt,~\psi_{0}\in\mathcal{D}(A).$ (18)
 $\rho(A)$ will contain a set of the form $\{s\ \in\mathbb{C};~{\rm Re}(s)\geq\omega\}$, where $\omega$ defines a growth bound on the semigroup $\{T_{A}(t)\}$.
-When this semigroup is exponentially stable, one can take $\omega=0$ [14, Ch.
+When this semigroup is exponentially stable, one can take $\omega=0$ [7, Ch.
 2].
-Then the operator-valued transfer function representation of $\Sigma(A,B,C,D)$ is given by
+Then the operator-valued transfer function representation of a system (5) $\Sigma(A,B,C,D)$ is given by
 $C(sI-A)^{-1}B+D:\mathcal{D}(A)\rightarrow L^{2}(\mathbb{T}),~~s\in\rho(A).$ (19)
-We say this transfer function is strictly proper if $\lim_{s\rightarrow\infty}\left(C(sI-A)^{-1}B+D\right)=0,$ i.e. if $D=0$.
-When $\Sigma(A,B,C,D)$ is spatially invariant, each of $A,B,C,D$ can be transformed to a multiplication operator on $\ell^{2}(\mathbb{Z})$ spaces of appropriate dimensions, e.g., $M_{\hat{A}}=\mathcal{F}^{-1}A\mathcal{F}$.
-In this case, (17) can be represented as
-$\tfrac{d}{dt}\hat{{\boldsymbol{\psi}}}(t)$ $=(M_{\hat{A}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{B}}\hat{\bf u})(t)$ (13)
-$\hat{\bf z}(t)$ $=(M_{\hat{C}}\hat{{\boldsymbol{\psi}}})(t)+(M_{\hat{D}}\hat{\bf u})(t),$
-and its transfer function can also be represented as a multiplication operator [14, Ch.7]:
-$\mathcal{F}^{-1}\left(C(sI-A)^{-1}B+D\right)\mathcal{F}=M_{\hat{G}(s)}$ (14)
-$\hat{G}_{\kappa}(s)=\hat{C}_{\kappa}(sI-\hat{A}_{\kappa})^{-1}\hat{B}_{\kappa}+\hat{D}_{\kappa}.$
-III-A Controller Design for the Diffusion Equation
-We focus here on a specific plant - the diffusion equation - noting that our approach could be generalized to other linear distributed parameter systems.
-The diffusion equation over $\mathbb{T}$ with fully distributed actuation $u$ and disturbance $w$ is described by the partial differential equation (PDE)
-$\partial_{t}\psi(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)+u(t,\theta)+w(t,\theta),~t\in[0,\infty),~\theta\in\mathbb{T},$ (15)
-where $\alpha>0$ is the constant of diffusivity. The spatio-temporal signal
-${z}(t,\theta)=\left[\begin{array}[]{c}\gamma{\psi}(t,\theta)\\ {u}(t,\theta)\end{array}\right],$ (16)
-specifies a performance output of interest that provides a uniform penalty across the spatial variable, with $\gamma>0$ determining a tradeoff between cost of state and cost of control.
-The dynamics (15)-(16) can be written as
-$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+{\bf u}(t)+{\bf w}(t)$ (17)
-${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
-where $A\in\mathcal{L}(L_{1}^{2}(\mathbb{T}))$ is defined by
-$(A\psi)(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)$ (18)
-and generates a $C_{0}$- semigroup $\{T_{A}(t)\}$ on
-$\{f\in L^{2}(\mathbb{T});~f^{\prime},f^{\prime\prime}\in L^{2}(\mathbb{\mathbb{T}})\}=:\mathcal{D}(A)\subset L^{2}(\mathbb{T}),$ (19)
-with derivatives, $f^{\prime},f^{\prime\prime}$, in (19) interpreted weakly, and $C,D\in\mathcal{L}(L_{1}^{2}(\mathbb{T}),L_{2}^{2}(\mathbb{T}))$ are the bounded operators
-$(C\psi)(t,\theta)=\left[\begin{array}[]{c}\gamma\psi(t,\theta)\\ 0\end{array}\right],~(Du)(t,\theta)=\left[\begin{array}[]{c}0\\ u(t,\theta)\end{array}\right].$ (20)
-We denote the operator-valued transfer function for (17) by
-$P(s)=C(sI-A)^{-1}+D,$ (21)
-which is well-defined on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\subset\rho(A)$ since 0 defines a growth bound on the semigroup $\{T_{A}(t)\}.$ $\Sigma(A,I,C,D)$ is a spatially-invariant system that is equivalently characterized by its transformation under $\mathcal{F}$ as
-$\tfrac{d}{dt}\hat{\psi}(t,\kappa)=\underbrace{-\alpha\kappa^{2}}_{\hat{A}_{\kappa}}\hat{\psi}(t,\kappa)+\hat{u}(t,\kappa)+\hat{w}(t,\kappa),$ (22)
-$\hat{z}(t,\kappa)=\underbrace{\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right]}_{\hat{C}_{\kappa}}\hat{\psi}(t,\kappa)+\underbrace{\left[\begin{array}[]{c}0\\ 1\end{array}\right]}_{\hat{D}_{\kappa}}\hat{u}(t,\kappa),~~t\in[0,\infty),~\kappa\in\mathbb{Z}$
-or in transfer function form as $\mathcal{F}^{-1}P(s)\mathcal{F}=M_{\hat{P}(s)},$
-$\hat{P}_{\kappa}(s)=\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right](s+\alpha\kappa^{2})^{-1}+\left[\begin{array}[]{c}0\\ 1\end{array}\right].$ (23)
-We consider a state feedback control policy defined by a transfer function
+If $A$ generates an exponentially stable $C_{0}$-semigroup, then we say that the transfer function representation of $\Sigma(A,B,C,D)$ is exponentially stable.
+If $D\equiv 0$, the transfer function representation of $\Sigma(A,B,C,D)$ is strictly proper.
+Now let $\Sigma(A,B,C,D)$ describe the diffusion equation (13).
+We denote the operator-valued transfer function from inputs $\begin{bmatrix}\boldsymbol{w}&\boldsymbol{u}\end{bmatrix}^{\top}$ to $\begin{bmatrix}\boldsymbol{z}&\boldsymbol{\psi}\end{bmatrix}^{\top}$ for (13) by $P(s)$, which is well-defined on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\subset\rho(A)$ as 0 defines a growth bound on the semigroup $\{T_{A}(t)\}.$ Since (13) is a spatially invariant system, $P(s)$ can be expressed as a multiplication operator, $M_{\hat{P}(s)},$ over spatial frequency [7, Ch.7], and (LABEL:eq:open_loop_kappa) can be represented by the transfer function
+$\hat{P}_{\kappa}(s)=\begin{bmatrix}\hat{C}_{\kappa}(s+\alpha\kappa)^{-1}&\hat{C}_{k}(s+\alpha\kappa)^{-1}+\hat{D}_{\kappa}\\ (s+\alpha\kappa)^{-1}&(s+\alpha\kappa)^{-1}\end{bmatrix}$, (20)
+which maps $\begin{bmatrix}\hat{w}_{\kappa}(s)&\hat{u}_{\kappa}(s)\end{bmatrix}^{\top}$ to $\begin{bmatrix}\hat{z}_{\kappa}(s)&\hat{\psi}_{\kappa}(s)\end{bmatrix}^{\top}$.
+We consider a (possibly dynamic) state feedback control policy defined by a spatially invariant transfer function
 ${\bf U}(s)=K(s){\boldsymbol{\Psi}}(s),$ (21)
 defined on some set $\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$.
-The feedback interconnection of (17) and (24) is a well-defined transfer function on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\cap\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$ [14, Ex.
-7.3], which we denote by $\mathbb{F}(P;K)(s).$
-Definition 2
-We say that the control policy (24) internally stabilizes the system (17) if $\mathbb{F}(P;K)(s)$ is the operator-valued transfer function of a system that generates an exponentially stable $C_{0}$-semigroup $\{T_{A_{\rm c.l.}}(t)\}$ and $\|\mathbb{F}(P;K)\|_{H_{2}}<\infty$.
-We now state formally the $H_{2}$ design problem of interest:
-$\inf_{K}~~\|\mathbb{F}(P;K)\|_{H_{2}}$ (25)
-$~{\rm s.t.}~~K~\text{spatially-invariant}$
-$~~~~~~K~\text{internally stabilizing}.$
-As in the finite-dimensional setting, this $H_{2}$ problem will recover the solution to the Linear Quadratic Regulator (LQR) problem for the diffusion equation [15, 14]:
-$\inf_{u}~\int_{0}^{\infty}\left<{\boldsymbol{\psi}}(t),\gamma^{2}{\boldsymbol{\psi}}(t)\right>_{L^{2}(\mathbb{T})}+\left<{\bf u}(t),{\bf u}(t)\right>_{L^{2}(\mathbb{T})}dt$ (26)
-[…]
-IV Closed-Loop Formulation of $H_{2}$ Problem
-In this section, we demonstrate that the controller design problem (LABEL:eq:H2_formulation) can be formulated as an optimization problem over certain closed-loop mappings and admits analytical solutions in the spatial frequency domain.
-We will illustrate that this formulation allows for spatio-temporal constraints to be imposed in a convex manner in Section VII.
-We isolate the closed-loop transfer functions from $w$ to $\psi$ and $u$ in the interconnection of (17) and (24) to obtain:
+The feedback interconnection of $P(s)$ and (21), denoted $\mathbb{F}(P;K)(s)$, is a well-defined transfer function on $\{s\in\mathbb{C};{\rm Re}(s)>0\}\cap\{s\in\mathbb{C};~{\rm Re}(s)\geq\omega\}$ [7, Ex.
+7.3] that maps $\boldsymbol{w}(s)$ to $\boldsymbol{z}(s)$.
+IV-B Closed-Loop Mappings
+We isolate the closed-loop transfer functions from $\boldsymbol{w}$} to $\boldsymbol{\psi}$} and $\boldsymbol{u}$} in $\mathbb{F}(P;K)(s)$ to obtain:
 ${\boldsymbol{\Psi}}(s)$ $=(I-(sI-A)^{-1}K(s))^{-1}(sI-A)^{-1}{\bf W}(s)$ (22)
@@
 $=:\Phi^{\psi}(s){\bf W}(s)$
-${\bf U}(s)$ $=K(s)(I-(sI-A)^{-1}K(s))^{-1}(sI-A)^{-1}{\bf W}(s)$
-$=K(s)(sI-A-K(s))^{-1}{\bf W}(s)$
+${\bf U}(s)$ $=K(s)(sI-A-K(s))^{-1}{\bf W}(s)$
 $=:\Phi^{u}(s){\bf W}(s).$
-The operator-valued transfer functions $\Phi^{\psi}(s),\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ are analogous to the “system-level maps” introduced for finite dimensional discrete-time systems in [5].
-The closed-loop transfer function from $w$ to $z$ can be written in terms of $\Phi^{\psi}(s),\Phi^{u}(s)$ as
+The operator-valued transfer functions $\Phi^{\psi}(s),\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ for each $s\in\mathbb{C}$ at which they are defined and are analogous to the “system-level maps” introduced for finite dimensional systems in [22].
+The closed-loop transfer function from $\boldsymbol{w}$} to $\boldsymbol{z}$} can be written in terms of $\Phi^{\psi}(s),\Phi^{u}(s)$ as
 ${\bf Z}(s)$ $=\mathbb{F}(P;K)(s){\bf W}(s)$ (23)
@@
 Proof:
-See Appendix -C. ∎
-Given the analogy to the definition for finite dimensional systems, we say that an operator-valued transfer function is an element of $H_{\infty}$ if it satisfies the condition (LABEL:eq:Hinf_condition).
+See Appendix -B. ∎
+Given the analogy to the definition for finite dimensional systems, we say that an operator-valued transfer function is an element of $\mathcal{H}_{\infty}$ if it satisfies the condition (LABEL:eq:Hinf_condition).
 Lemma 2
-Let $\Phi^{\psi}(s)$ and $\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ be defined by (27) for some spatially-invariant and internally stabilizing controller $K(s).$ Then, for all $\eta,\vartheta\in L^{2}(\mathbb{T})$ and $s\in\{s\in\mathbb{C};{\rm Re}(s)\geq 0\}$,
+Let $\Phi^{\psi}(s)$, $\Phi^{u}(s)\in\mathcal{L}(L^{2}(\mathbb{T}))$ be defined by (22) for some spatially-invariant and internally stabilizing controller $K(s).$ Then, for all $\eta,\vartheta\in L^{2}(\mathbb{T})$ and $s\in\{s\in\mathbb{C};{\rm Re}(s)\geq 0\}$,
 $\left<\eta,\left((s+1)I-(A+I)\right)\Phi^{\psi}(s)\vartheta\right>_{L^{2}(\mathbb{T})}$ (26)
@@
 Proof:
-See Appendix -D. ∎
+See Appendix -C. ∎
+We can now state our main result:
+Theorem 1
+The solution to the $\mathcal{H}_{2}$ design problem for the diffusion equation (17) can be recovered from the solution to:
+$\operatorname*{arg\,min}_{\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{\mathcal{H}_{2}}$ (28)
+$~~~~~{\rm s.t.}~~~~~~\Phi^{\psi},\Phi^{u}\text{ spatially invariant, strictly proper}$
+[…]
+Lemmas 1-2 show that $K$ spatially invariant and internally stabilizing implies that $\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}$ and are spatially invariant, strictly proper, and satisfy the affine constraint (LABEL:eq:affine_innerProduct).
+To complete the proof of Theorem 1 it remains to show that any such $\Phi^{\psi},\Phi^{u}$ can be implemented as an internally stabilizing controller.
+We will prove this in Lemma 3 and discuss controller implementation in Section V.
+IV-C Analytical Solution to Closed-Loop Formulation of $\mathcal{H}_{2}$ Problem
+Spatial invariance allow us to solve (LABEL:eq:CL_formulation) in the spatial frequency domain.
+By the Plancherel theorem, formulation (LABEL:eq:CL_formulation) of the $\mathcal{H}_{2}$ design problem for the diffusion equation is equivalent to the optimization problem:
+$\operatorname*{arg\,min}_{{\hat{\Phi}^{\psi}(s)},{\hat{\Phi}^{u}(s)}}~\left\|M_{\hat{C}}M_{\hat{\Phi}^{\psi}(s)}+M_{\hat{D}}M_{\hat{\Phi}^{u}(s)}\right\|_{\mathcal{H}_{2}}$ (29)
+$~~~~~{\rm s.t.}~~~~~~\hat{\Phi}^{\psi},\hat{\Phi}^{u}\text{ strictly~proper,~spatially~invariant},$
+[…]
+Note that the cost and constraints of (LABEL:eq:CLOpt_Frequency) “decouple” in the spatial frequency variable, so the solution to (LABEL:eq:CLOpt_Frequency) can be obtained by solving a parameterized family of $\mathcal{H}_{2}$ problems for finite-dimensional systems parameterized by $\kappa\in\mathbb{Z}$.
+Let ${\mathcal{RH}}_{2}$ denote the set of rational transfer functions with strictly proper entries and no poles in the closed right half plane [8, Ch.
+3].
+Then the symbol of the minimizer of (LABEL:eq:CLOpt_Frequency) at frequency $\kappa$ is the solution to the family of problems:
+$\operatorname*{arg\,min}_{\hat{\Phi}^{\psi}_{\kappa},\hat{\Phi}^{u}_{\kappa}\in\mathcal{RH}_{2}}\left\|\left[\begin{array}[]{cc}\gamma&0\\ 0&1\end{array}\right]\left[\begin{array}[]{c}\hat{\Phi}^{\psi}_{\kappa}\\ \hat{\Phi}^{u}_{\kappa}\end{array}\right]\right\|_{\mathcal{H}_{2}}$ (30e)
+$~~{\rm s.t.}\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,$ (30f)
+provided that the family of solutions $\{\hat{\Phi}^{\psi}_{\kappa}\},~\{\hat{\Phi}^{u}_{\kappa}\}$ satisfy the condition (LABEL:eq:Hinf_frequency) which ensures $\Phi^{\psi},\Phi^{u}\in\mathcal{H}_{\infty}$.
+For each fixed $\kappa\in\mathbb{Z},$ (30) is a standard $\mathcal{H}_{2}$ design problem for a finite-dimensional system with an implicit affine subspace constraint (30f), which we can solve analytically.
+Proposition 1
+The optimal solution $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ to the problem (30), with $\gamma$ defined in (12), is given by
+$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\left(s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}\right)^{-1},$ (31)
+$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}.$
+Proof:
+See Appendix -A. ∎
+V internally stable implementation of Controller via Closed-Loop Solution
+We now complete the proof of Theorem 1 by showing that $\Phi^{\psi},\Phi^{u}$ satisfying the constraints of (LABEL:eq:CL_formulation) can be used to construct an internally stable implementation of the controller, which thus preserves the structural properties of these maps.
+In order for an implementation of the controller to be well posed and preserve the structure of $\Phi^{\psi}$ and $\Phi^{u}$, we introduce an additional transfer function $G$ that ensures internal states of the controller remain bounded.
 Lemma 3
-Let ${\Phi}^{\psi},{\Phi}^{u}\in H_{\infty}\subset\mathcal{L}(L^{2}(\mathbb{T}))$ be spatially-invariant, strictly proper, and satisfy (32).
-Select a spatially-invariant transfer function [^1] Existence of such a $G$ will be verified in Section V-A.
-$G$ for which $G^{-1}(s)\in H_{\infty}$, $\lim_{s\to\infty}s\hat{G}_{\kappa}^{-1}(s)=1~\forall\kappa\in\mathbb{Z}$, and $G^{-1}(s)(sI-A)\in H_{\infty}$.
+Let ${\Phi}^{\psi},{\Phi}^{u}\in\mathcal{H}_{\infty}\subset\mathcal{L}(L^{2}(\mathbb{T}))$ be spatially-invariant, strictly proper, and satisfy (27).
+Select a spatially-invariant transfer function $G$ for which $G^{-1}(s)\in\mathcal{H}_{\infty}$, $\lim_{s\to\infty}s\hat{G}_{\kappa}^{-1}(s)=1~\forall\kappa\in\mathbb{Z}$, and $G^{-1}(s)(sI-A)\in\mathcal{H}_{\infty}$.
 Then,
@@
 ${\bf U}(s)$ $=G(s){\Phi}^{u}(s){\bf V}(s)$
-defines an internally stabilizing controller for (17).
-This implementation (33) is well-posed and the internal signal $\bf v$} remains bounded for all bounded signals $\bf n$} and $\bf w$} as depicted in Figure 1.
+defines an internally stable implementation of the controller in the sense that the closed loop transfer matrices from the exogenous inputs $(\hat{n}_{\kappa}(s),\hat{w}_{\kappa}(s))$ to the internal signals $(\hat{\psi}_{\kappa}(s),\hat{u}_{\kappa}(s),\hat{v}_{\kappa}(s))$ belong to $\mathcal{H}_{\infty}$.
+This implementation is depicted in Figure 1.
 The input-output map (transfer function) corresponding to (32) is
-${\bf U}=\mathcal{F}^{-1}M_{\hat{K}(s)}\mathcal{F}{{\boldsymbol{\Psi}}},~~\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}.$ (34)
+${\bf U}=\mathcal{F}^{-1}M_{\hat{K}(s)}\mathcal{F}{{\boldsymbol{\Psi}}},~~\hat{K}_{\kappa}(s)=\tfrac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}.$ (33)
 Proof:
-See Appendix -E. ∎
+See Appendix -D. ∎
 Remark 2
@@
 Figure 1: Block diagram of the dynamic controller implementation (32) with noisy measurements and exogenous disturbance: $w$ is additive noise to the state $\psi$, and $w$ is additive noise to the control signal $u$.
-Lemmas 1-3 demonstrate that the set of all internally stabilizing and spatially-invariant controllers for (17) can be characterized by an affine subspace of spatially-invariant transfer functions $\Phi^{\psi},\Phi^{u}.$ With this, we are able to state our main result.
-Theorem 1
-The following is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation) for the diffusion equation:
-$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (35)
-$~~~~~{\rm s.t.}~~~~~~\Phi^{\psi},\Phi^{u}\text{ spatially invariant, strictly proper}$
-[…]
-V Analytical Solution to Closed-Loop Formulation of $H_{2}$ Problem
-Spatial invariance properties allow us to solve (LABEL:eq:CL_formulation) in the spatial frequency domain.
-By the Placherel theorem, formulation (LABEL:eq:CL_formulation) of the $H_{2}$ design problem for the diffusion equation is equivalent to the optimization problem:
-$\inf_{{\hat{\Phi}^{\psi}(s)},{\hat{\Phi}^{u}(s)}}~\left\|M_{\hat{C}}M_{\hat{\Phi}^{\psi}(s)}+M_{\hat{D}}M_{\hat{\Phi}^{u}(s)}\right\|_{H_{2}}$ (36)
-$~~~~~{\rm s.t.}~~~~~~\hat{\Phi}^{\psi},\hat{\Phi}^{u}\text{ strictly~proper,~spatially~invariant},$
-[…]
-Note that the cost and constraints of (LABEL:eq:CLOpt_Frequency) “decouple” in the spatial frequency variable $\kappa.$ Thus, the solution to (LABEL:eq:CLOpt_Frequency) can be obtained by solving the family of $H_{2}$ problems, parameterized by $\kappa\in\mathbb{Z}:$
-$\inf_{\hat{\Phi}^{\psi}_{\kappa},\hat{\Phi}^{u}_{\kappa}}\left\|\left[\begin{array}[]{cc}\gamma&0\\ 0&1\end{array}\right]\left[\begin{array}[]{c}\hat{\Phi}^{\psi}_{\kappa}\\ \hat{\Phi}^{u}_{\kappa}\end{array}\right]\right\|_{H_{2}}$ (37e)
-$~~{\rm s.t.}~~~\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ (37f)
-$~~~~~~~~\left((s+1)-(-\alpha\kappa^{2}+1)\right)\hat{\Phi}^{\psi}_{\kappa}(s)-\hat{\Phi}_{\kappa}^{u}(s)=1,$ (37g)
-provided that the family of solutions $\{\hat{\Phi}^{\psi}_{\kappa}\},~\{\hat{\Phi}^{u}_{\kappa}\}$ satisfy the condition (LABEL:eq:Hinf_frequency) which ensures $\Phi^{\psi},\Phi^{u}\in H_{\infty}$.
-For each fixed $\kappa\in\mathbb{Z},$ (37) is a standard $H_{2}$ design problem for a finite-dimensional system with an implicit affine subspace constraint (37g).
-To obtain an analytical solution to this problem, we first replace the implicit constraint (37g) with an analogous explicit constraint.
-Lemma 4
-Fix $\kappa\in\mathbb{Z}.$ The single-input single-output (SISO) transfer functions $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ satisfy $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}\in RH_{2}$ and (37g) if and only if they are of the form
-$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+1}\left(1+\hat{\rho}_{\kappa}(s)\right),$ (38)
-$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-1}{s+1}+\frac{s+\alpha\kappa^{2}}{s+1}\hat{\rho}_{\kappa}(s)$
-for some SISO transfer function $\hat{\rho}_{\kappa}(s)\in RH_{2}.$
-Proof:
-See Appendix -F. ∎
-We leverage the parameterization (38) and properties of the $H_{2}$ norm to write (37) as an unconstrained $H_{2}$ design problem over a single SISO transfer function decision variable $\hat{\rho}_{\kappa}.$
-Proposition 1
-The optimal solution $\hat{\Phi}^{\psi}_{\kappa},~\hat{\Phi}^{u}_{\kappa}$ to the problem (37) can be recovered from the solution $\hat{\rho}_{\kappa}$ to the optimization problem
-$\inf_{\hat{\rho}_{\kappa}\in RH_{2}}~\left\|\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{\alpha\kappa^{2}-1}{s+1}\end{array}\right]+\left[\begin{array}[]{c}\tfrac{\gamma}{s+1}\\ \tfrac{s+\alpha\kappa^{2}}{s+1}\end{array}\right]\hat{\rho}_{\kappa}(s)\right\|_{H_{2}}$ (39)
-through the relation (38). The optimal solution to (39) is
-$\hat{\rho}_{\kappa}^{\rm opt.}(s)=\frac{1-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (40)
-and thus the optimal solution to (37) is given by
-$\hat{\Phi}^{\psi}_{\kappa}(s)$ $=\frac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}},$ (41)
-$\hat{\Phi}^{u}_{\kappa}(s)$ $=\frac{\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}.$
-Proof:
-See Appendix -A. ∎
-V-A Implementation of Controller via Closed-Loop Solution
 We examine the structure of the controller implementation (32) that results from the optimal solution (31). We select $G(s)=\mathcal{F}^{-1}M_{\hat{G}(s)}\mathcal{F}$ where
 $\hat{G}_{\kappa}(s)=s+1+\kappa^{2},$ (34)
-to satisfy all conditions of Lemma 3. In this case, (33) may be specified in spatial frequency $\kappa$ as
+to satisfy all conditions of Lemma 3. Note that this is one suitable choice of $G$, but it is not obvious how to “best” select $G$ in general. Then, (32) is specified at frequency $\kappa$ as
 $\hat{v}(s,\kappa)$ $=\hat{\psi}(s,\kappa)+\left(1-\hat{G}_{\kappa}(s)\tfrac{1}{s+\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}}\right)\hat{v}(s,\kappa)$ (35)
@@
 $\begin{bmatrix}\hat{\psi}_{\kappa}\\ \hat{u}_{\kappa}\\ \hat{v}_{\kappa}\end{bmatrix}=\begin{bmatrix}\hat{\Phi}^{u}_{\kappa}(s)&\hat{\Phi}^{\psi}_{\kappa}(s)\\ (s+\alpha\kappa^{2})\hat{\Phi}^{u}_{\kappa}(s)&\hat{\Phi}^{u}_{\kappa}(s)\\ (s+\alpha\kappa^{2})\hat{G}^{-1}_{\kappa}(s)&\hat{G}^{-1}_{\kappa}(s)\end{bmatrix}\begin{bmatrix}\hat{n}_{\kappa}\\ \hat{w}_{\kappa}\\ \end{bmatrix}$, (36)
-each of which is verified to be the symbol of a multiplication operator in $H_{\infty}$ for our selected $\hat{G}$.
+each of which is verified to be the symbol of a multiplication operator in $\mathcal{H}_{\infty}$ for our selected $\hat{G}$.
 We numerically compute the spatio-temporal impulse responses (Green’s functions) of the blocks in Fig.
-1 and plot their dynamic and static components in Fig.
+1 and plot their dynamic and static components over the spatial domain for various selections of time in Fig.
 2.
@@
 VI Equivalence to LQR Control
-We demonstrate that the solution to our $H_{2}$ design problem indeed recovers the solution to the standard infinite-horizon LQR problem for the diffusion equation.
+We demonstrate that the solution to our closed-loop $\mathcal{H}_{2}$ design problem (LABEL:eq:CL_formulation) recovers the solution to the standard infinite-horizon LQR problem for the diffusion equation.
 Proposition 2
-Let $(\mathcal{F}\psi_{0})(\kappa)=1$ for all $\kappa$.
-[^2] Note that the “response” (weak solution) of a system to this initial condition is equal to the Green’s Function of the linear PDE describing the system dynamics, and that $\mathcal{F}$ here is the extension of the Fourier transform to the space of distributions.
-Then the standard LQR problem (LABEL:eq:LQR) is equivalent to the $H_{2}$ problem of interest (LABEL:eq:H2_formulation).
+The standard LQR problem (LABEL:eq:LQR) and the $\mathcal{H}_{2}$ problem of interest (LABEL:eq:CL_formulation) have the same solution $K^{\rm opt}$.
 Proof:
-First, we compute the solution to the LQR formulation (LABEL:eq:LQR) using standard approaches [16, 14]. We solve the operator-valued Ricatti equation
+First, we compute the solution to the LQR formulation (LABEL:eq:LQR) using standard approaches [2, 7]. We solve the operator-valued Riccati equation
 $\left<A\eta,F\vartheta\right>_{L^{2}(\mathbb{T})}+\left<F\eta,A\vartheta\right>_{L^{2}(\mathbb{T})}+\gamma^{2}\left<\eta,\vartheta\right>_{L^{2}(\mathbb{T})}$ (37)
@@
 $\alpha\kappa^{2}\hat{F}_{\kappa}-\hat{F}_{\kappa}\alpha\kappa^{2}-\hat{F}_{\kappa}^{2}+\gamma^{2}=0.$ (38)
-The unique solution $\hat{F}_{\kappa}>0$ to (46) is $\hat{F}_{\kappa}=-\alpha\kappa^{2}+\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}},$ and the optimal state feedback gain is thus given by
-$K=\mathcal{F}M_{\hat{K}}\mathcal{F}^{-1},~~\hat{K}_{\kappa}=-\hat{F}_{\kappa}=\alpha\kappa^{2}-\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}}.$ (47)
+The unique solution $\hat{F}_{\kappa}>0$ to (38) is $\hat{F}_{\kappa}=-\alpha\kappa^{2}+\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}},$ and the optimal state feedback gain is thus given by $K^{\rm LQR}=\mathcal{F}M_{\hat{K}^{\rm LQR}}\mathcal{F}^{-1}$ where
+$\hat{K}^{\rm LQR}_{\kappa}=-\hat{F}_{\kappa}=\alpha\kappa^{2}-\sqrt{\alpha^{2}\kappa^{4}+\gamma^{2}}.$ (39)
 We next obtain the input-output mapping specified by the controller implementation (35) using (33) as
-$\hat{K}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}=\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}.$ (48)
-This is indeed equal to the static state feedback gain $\hat{K}$ obtained from the Riccati equation solution.
-Details on the equivalence of the cost of the two formulations (LABEL:eq:CL_formulation) and (LABEL:eq:LQR) are provided in Appendix -B.
-∎
-We know that $K$ will take the form of a spatial convolution operator
+$\hat{K}^{\mathcal{H}_{2}}_{\kappa}(s)=\frac{\hat{\Phi}^{u}_{\kappa}(s)}{\hat{\Phi}^{\psi}_{\kappa}(s)}=\alpha\kappa^{2}-\sqrt{\gamma^{2}+\alpha^{2}\kappa^{4}}.$ (40)
+This is indeed equal to the static state feedback gain $\hat{K}^{LQR}$, thus $K^{\rm LQR}=K^{\mathcal{H}_{2}}=K^{\rm opt}.$
+We know that $K^{\rm opt}$ will be a spatial convolution operator
 $u(t,\theta)=\int_{\xi\in\mathbb{T}}k(\theta-\xi)\psi(t,\xi)dt,$ (41)
 where $k(\theta)=(\mathcal{F}^{-1}\hat{K})(\theta).$ Numerical integration shows that $k(\theta)$ exhibits spatial decay away from the origin (see Fig. 2).
-We compare the static controller $K$ to the dynamic controller implementation (43) depicted in Fig.
+We compare the static controller $K^{\rm opt}$ to the dynamic controller implementation (35) depicted in Fig.
 1.
@@
 The transfer function $G(s)\Phi^{u}(s)$ has a corresponding convolution kernel with both a static and dynamic part, which we denote as $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta)$, respectively.
-The static (feedthrough) component of $G(s)\Phi^{u}(s)$ is obtained by evaluating its Fourier symbol in the limit as $s\to\infty$,
-$\mathcal{F}^{-1}\lim_{s\to\infty}\hat{G}(s,\kappa)\hat{\Phi}^{u}(s,\kappa)=h_{\Phi^{u}}(\theta)=(\mathcal{F}^{-1}\hat{K})(\theta),$
+The static (direct feedthrough) component of $G(s)\Phi^{u}(s)$ is obtained by evaluating its Fourier symbol in the limit as $s\to\infty$,
+$\mathcal{F}^{-1}\lim_{s\to\infty}\hat{G}(s,\kappa)\hat{\Phi}^{u}(s,\kappa)=h_{\Phi^{u}}(\theta)=(\mathcal{F}^{-1}\hat{K}^{\rm opt})(\theta),$
 which shows that the static part of the implementation (35) recovers the static LQR gain, as expected.
-The “dynamic components” of the systems $G(s)\Phi^{u}(s)$ and $\left(I-G(s){\Phi}^{\psi}(s)\right)$ cancel each other out, leaving only the static feedback policy $K.$ This is confirmed in the plot of $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta),f_{\Phi^{\psi}}(t,\theta)$ shown in Fig.
+The “dynamic components” of the systems $G(s)\Phi^{u}(s)$ and $\left(I-G(s){\Phi}^{\psi}(s)\right)$ cancel each other out, leaving only the static feedback policy $K^{\rm opt}.$ This is confirmed in the plot of $h_{\Phi^{u}}(\theta),f_{\Phi^{u}}(t,\theta),f_{\Phi^{\psi}}(t,\theta)$ shown in Fig.
 2.
 VII Incorporation of Spatio-Temporal Decay Rate Constraints
-Without constraints, the closed-loop formulation of the $H_{2}$ problem (LABEL:eq:CL_formulation) is equivalent to a standard LQR problem.
+Without constraints, the closed-loop formulation of the $\mathcal{H}_{2}$ problem (LABEL:eq:CL_formulation) is equivalent to a standard LQR problem.
 The advantage of this closed-loop formulation is that it remains convex when spatio-temporal constraints are incorporated.
@@
 An added constraint that $I-G\Phi^{\psi},G\Phi^{u}\in\mathcal{S}$ enforces a structural constraint on the components of our controller. In this case, formulation (LABEL:eq:CL_formulation) becomes:
-$\inf_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{H_{2}}$ (50)
+$\operatorname*{arg\,min}_{\Phi^{\psi},\Phi^{u}\in H_{\infty}}~\left\|C\Phi^{\psi}+D\Phi^{u}\right\|_{\mathcal{H}_{2}}$ (42)
 $~~~~~{\rm s.t.}~~~~~~\Phi^{\psi},\Phi^{u}\text{ spatially invariant}$
@@
 One possible spatio-temporal constraint is when $\mathcal{S}$ is the set of linear PDEs whose Green’s functions $\rho(\cdot,\cdot)$ satisfy
-$\|\rho(\cdot,t)-\rho_{T}(\cdot,t)\|_{L_{2}(\mathbb{T})}<\epsilon(t)$
+$\|\rho(\cdot,t)-\rho_{T}(\cdot,t)\|_{L_{2}(\mathbb{T})}<\epsilon(t)$ (43)
 where $\rho_{T}$ is the spatial truncation of $\rho$ given by
@@
 Choosing $\epsilon(t)$ to grow in $t$ allows $\rho$ to be less localized in space as time increases.
+While constraining $\rho$ to have compact support as in [5] admits desirable computational properties, the decay constraint (43) is perhaps more physically relevant.
 This set $\mathcal{S}$ is convex, and guarantees that a spatially truncated version of controller will perform well using small gain arguments.
@@
 VIII Conclusion
-It was shown that the infinite-horizon LQR problem for the diffusion equation over the unit circle can be reformulated as an $H_{2}$ problem over the system’s closed-loop responses, analogous to the finite-dimensional System Level Synthesis methodology.
-Implementation of the resulting controller using the closed-loop responses was shown to preserve stability of the closed-loop system.
-It was further illustrated that this novel formulation admits constraints on the closed-loop responses while preserving convexity.
-Generalizations of the presented results to a class of linear distributed parameter systems over both compact and unbounded spatial domains are the subject of ongoing work.
-Of particular interest is comparing analytical results from this framework to the computational approach of [9].
+It was shown that the infinite-horizon LQR problem for the diffusion equation over the unit circle can be reformulated as an $\mathcal{H}_{2}$ problem over the system’s closed-loop responses, analogous to the finite-dimensional System Level Synthesis methodology.
+An implementation of this optimal controller built from these closed-loop responses was shown to preserve internal stability.
+It was further illustrated that this closed-loop formulation admits a class of spatio-temporal constraints on the closed-loop responses while preserving convexity.
+Generalizations of this work to general linear distributed parameter systems are the subject of ongoing work.
+Of particular interest include incorporation of boundary conditions through an embedding approach [9] or alternate representations [17]; and comparison of analytical results from this work to the computational approach of [6].